OpenStudy (anonymous):

A bread recipe calls for 1 1/2 cups of regular flour and 1 1/5 cups of wheat flour. What is the total amount flour is needed?

OpenStudy (anonymous):

1 1/2 =3/2 1 1/5=6/5 total=3/2+6/5=(15+12)/10=27/10